if vectors 2i-3j+k, 3i+xj+k are orthogonal then x=?

Grade:9

3 Answers

KARNEYSANKETH
7 years ago
If two vectors are orthogonal means that they are in perpendicular to each other and the dot product is zero. here, dot product of two vectors is 6-3x+1=5-3x=0...x=5/3..It`s m
Kuldeep Pal
7 years ago
Orthogonal vectors means that dot product is Zero and hence:
(2)(3) + (3)(x) + (k)(k) =0........(1)
 and for 2nd equation ,
we have Cross product as Zero.
from Both equation we can find ‘k’ and ‘x’ 
kishan ravishankar
6 years ago
If the 2 vectors are orthogonal, they are said to be perpendicular to each other.
Therefore, their dot product is equal to 0.
(2*3)+(-3*x)+(1*1)=0         {Since i.i=1 , j.j=1 , k.k=1}
6-3x+1=0
3x=7
x=7/3=2.3